An inductor-coil, a capacitor and an AC source of rms voltage 24 V are connected in series. When the frequency of the source is varied, a maximum rms current of 6.0 A is observed. If this inductor coil is connected to a battery of emf 12 V and interna resistance 4.0 Ω, what will be the current?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. E base rms = 24 V R = 4 Ω, I base rms = 6 A R = E/I = 24/6 = 4 Ω Internal Resistance = 4 Ω Hence net resistance = 4 + 4 = 8 Ω ∴ Current = 12/8 = 1.5 A